A coil of copper wire has a resistance of 90ohm at 20°c and is conducted to a 230v supply. By how much should the voltage be increased to keep the current constant if the temperature of coil rises to 60°c


1
Expert's answer
2021-05-12T16:42:18-0400

Take the temperature coefficient of resistance of copper as 0.00428 from 0°C.

"\frac{R_{60}}{R_{20}}= \frac{1+60×0.00428}{1+20×0.00428}"


"R_{60°C.} = 90 × \frac {1.2568}{1.0856 }= 104.2Ω"

"R_{20°C }= \frac{230}{90} = 2.56A"


Since the wire resistance has become 104.2Ω at 60°C

 "New voltage= 104.2 \u00d7 2.56= 266.3 V"

∴ increase in voltage required = 266.3 − 230 = 36.3 VR